When the central statics office surveys the population ,it places all adults 16years and above into one of the category employed , unemployed or not in the labour force . zambia's population was estimated to be 17.4 million people as of 2018. When officiating at the pensions and insurance authority event. The minister of labour and social services in late 2018 disclosed that zambia's unemployment rate is 41.2. the working age population is 16 years or older was estimated at 9056840.The labour force population was 3398294. Calculate the total number of people unemployed ,the labour participation rate ?

1
Expert's answer
2021-12-17T11:58:02-0500

Unemployment rate ="\frac{Number of unemployed}{Labour force population}" "*" 100"\%" "%"

Number of unemployed= Unemployment rate "*" Labour force population

No. of unemployed = 41.2"\%" "*"3398294

= 1,400,098 People

Labour Participation rate

= "\frac{Labour force population}{Working age population}" "*" 100"\%"

= "\frac{3398294}{9056840}" "*" 100"\%"

= 37.523"\%"